OpenStudy (anonymous):

Given the point (5, 6) and the slope of 5, find y when x = 21. 86 90 105 111

OpenStudy (agreene):

start by putting it in slope point notation: \[y-y_1=m(x-x_1)\] \[y-6=5(x-5)\] then solve for y and then plug in the value for x

OpenStudy (agreene):

actually i suppose you could just plug in for x now and just move that right side so its just y after